# Role of interaction anisotropy in polymer cononsolvency: insights from the Flory-Huggins-Potts framework

### v2

@@
Department of Chemical and Biological Engineering, Princeton University, Princeton, NJ 08544.
Cononsolvency occurs when mixing two good solvents creates poor-solvent conditions for polymers over specific composition ranges, causing macroscopic phase separation or microscopic chain collapse. Despite its technological and biophysical relevance, the connection between macroscopic and microscopic manifestations of cononsolvency remains unclear. A key challenge is identifying which interactions govern cononsolvency: coarse-grained analyses like standard Flory–Huggins models assume purely isotropic interactions, while atomistic simulations contain complex anisotropic interactions that cannot be precisely controlled or isolated. Here, we address the role of interaction anisotropy using the Flory-Huggins-Potts framework, which treats $\chi$ as a thermodynamic average over molecular orientations. This enables controlled comparison between systems with isotropic versus anisotropic interactions that share identical effective $\chi$ parameters, either driving cononsolvency by strong solvent-cosolvent affinity or preferential polymer-cosolvent affinity. While pairs of systems exhibit equivalent macroscopic phase behavior, lattice Monte Carlo simulations reveal that those featuring anisotropic interactions generate distinct collapse signatures, particularly in reentrant coil-globule transitions or characteristics of the solvation structure. These results demonstrate how microscopic interaction anisotropy influences cononsolvency behavior beyond what effective $\chi$ parameters alone predict.
@@ 1 Introduction
Multiple mechanisms of polymer cononsolvency have been reported on the basis of various theoretical and computational analyses. The preferential polymer-cosolvent adsorption mechanism suggests that the cosolvent, on trace addition, selectively solvates the polymer or facilitates formation of “enthalpic bridges” between otherwise distal parts of a polymer chain.34 These interactions can then trigger phase separation or single-chain collapse.35, 36, 37, 38, 39, 40, 14, 41, 42 Effects akin to preferential adsorption can also be realized based purely on entropic factors, such as size asymmetry between solvent and cosolvent leading to effective depletion interactions.43 In another scenario, the cosolvent is envisioned to act like a surfactant, such that polymer contraction allows the cosolvent to interact favorably with both solvent and polymer.44 Another manifestation arises when solvent-cosolvent interactions are substantially more favorable than interactions with the polymer, such that the polymer sacrifices its entropy to encourage solvent-cosolvent mixing.13, 4, 16, 45, 46, 43 The concept of geometric frustration has also been used to explain cononsolvency, proposing that competition between cosolvent and solvent with polymer destabilizes the polymer solvation environment.47, 48 Overall, this literature highlights polymer cononsolvency as a complex phenomenon with various enthalpic and entropic factors informing distinct mechanistic views.
@@
Previous studies using Flory-Huggins (FH) theory have demonstrated that phase separation can be induced when either solvent-cosolvent or polymer-solvent interactions dominate, even when the polymer is miscible with each solvent individually.45, 49, 50 Furthermore, analyses of structure factors obtained following application of the random phase approximation provide similar conclusions. 51 Nevertheless, mean-field insights into the energetic factors driving phase separation do not directly translate into understanding microscopic physics, such as single-chain conformational behavior.34 In this direction, field theoretic methods14, 42 have demonstrated that single-chain collapse indeed occurs given strong, preferential affinity of the polymer for one of the solvents, but there is no apparent coil-globule transition driven purely by solvent-cosolvent affinity.
@@
Molecular dynamics (MD) simulations of systems with generic chemical attributes (e.g., bead-spring polymers in monomeric solvents) have effectively illustrated many of the aforementioned mechanisms of cononsolvency at a microscopic level.43, 34, 52 A typical observation from such simulations is that the $R_{\text{g}}$ for a single polymer chain first decreases and then increases as cosolvent is progressively added. This scenario has been induced by solvent-cosolvent mixing, preferential polymer-solvent adsorption, and the surfactant-like action of the cosolvent by analyzing the molecular environment of the polymer and carefully evaluating molecular affinities. 34, 53, 54, 55, 48 In chemically specific systems, such as poly(N-isopropylacrylamide) in water and ethanol, single-chain collapse, consistent with cononsolvency, has been observed, but there is debate as to what precise interactions beget this behavior.56, 57, 58, 59 While various forms of cononsolvency have been observed, mechanisms are typically studied using different models and approaches, with macroscopic (solution-phase) and microscopic (single-chain) behaviors rarely considered together.
@@
Several recent works have established quantitative connections between microscopic chain behavior and macroscopic phase separation in cononsolvency for systems with isotropic interactions. Zhang has comprehensively mapped the energetic conditions (i.e., $\chi$ parameters) that yield cononsolvency through mean-field analysis.50, 60 Previously, Zhang et al. also demonstrated via Wang-Landau simulations that trace cosolvent addition induces continuous coil-globule-coil transitions without requiring explicit solvent-solvent attraction.61 By a complmenetary and distinct approach, Marcato et al. mapped lattice models onto $\mathcal{O}(n)$-vector spin models to derive exact partition functions and field-theoretic descriptions amenable to analytics.62 Li et al. have established connections between single-chain collapse, multi-chain aggregation, and mean-field theory for block copolymers and homopolymer chain(s) in binary solvents, in the context of FH models63 While these studies establish that specific combinations of scalar $\chi$ parameters derived from isotoropic interactions produce cononsolvency, this leaves unresolved whether interaction anisotropy fundamentally alters cononsolvency mechanisms or merely modulates existing ones.
@@
While simple FH theory captures cononsolvency phenomena, the apparent connection between thermoresponsive polymer solutions and cononsolvency37, 64, 2 and highlighted importance of hydrogen bonding,65, 37 ultimately suggest that interaction anisotropy may be relevant for at least some specific cononsolvency mechanisms. Here, we directly investigate how interaction anisotropy influences polymer cononsolvency in the context of a Flory-Huggins-Potts (FHP) framework,66 which previously demonstrated that including orientation-dependent interactions within FH-like models enabled description of thermoresponsive phenomena (e.g., miscibility loops and heating-induced coil-globule transitions) without temperature-dependent $\chi$ parameters. We first establish consistency with established phase separation regimes driven by solvent-cosolvent mixing versus polymer-cosolvent adsorption. Then, we compare corresponding coil-globule-coil transitions across systems with identical effective $\chi$ parameters but different underlying energetic contributions using lattice Monte Carlo simulations. Subsequent analysis reveals how anisotropic interactions generate microscopically distinct collapse pathways compared to isotropic systems, despite yielding identical macroscopic $\chi$ parameters. These findings, which are obtained using a single theoretical framework, clarify how orientation-dependent interactions, like hydrogen bonding or even packing effects, may alter microscopic structuring in cononsolvency and suggest strategies for understanding or tuning polymer response based on molecular anisotropy.
@@ 2.1 Flory-Huggins-Potts framework and stability analysis
To enable description of orientation-dependent interactions, we extend the recently developed FHP framework66 to model ternary polymer solutions. In effect, the FHP framework augments conventional FH theory by assigning internal states to particles that modulate pairwise interactions. These states may be interpreted, for example, as capturing aspects of molecular orientation at a coarse-grained resolution, enabling expression of orientation-dependent interactions, whereas traditional FH assumes isotropic interactions. While the motivation and detailed development of the FHP framework can be found in Ref. 66 , we briefly review some salient aspects.
@@
where $\Theta(\cdot)$ is the Heaviside step function, and $\delta$ defines the angular tolerance for alignment. We refer to this interaction style as a “correlation network,” and it resembles interactions in Maier-Saupe theory.67 Correlation networks allow all neighbors to align cooperatively, since the function becomes nonzero when two vectors point nominally in the same direction. This is perhaps reminiscent to the collective ordering observed in nematic liquid crystals or hydrogen-bonding networks in water, where molecular orientations enable multiple simultaneous interactions.
We note that the transition from simple FH to FHP introduces additional parameters governing particle interactions. However, not all parameters are essential for modeling specific phenomena, and FHP parameters have been successfully fitted to experimental data with high precision. Regularization techniques and physical constraints can yield parsimonious models when needed. Since our objective is to distinguish cononsolvency manifestations, we employ minimal models that introduce only two new parameters per system while keeping other variables fixed.
@@ 2.2 Helmholtz energy by mean-field analysis
Following Ref.  66 , but extended to three components, the intensive free energy of mixing per lattice site is:
@@
The terms $p_{v}$ and $p_{\Omega}$ are geometric factors related to what defines aligned interactions. In particular, $p_{v}$ is the fraction of neighbors that can form aligned interactions, and $p_{\Omega}$ is the fraction of pairwise orientations that are classified as aligned for such neighbors.
For simplicity, all numerical results correspond to a simple cubic lattice where $\mathcal{N}(i)$ includes the nearest, next-nearest, and next-next-nearest neighbors of site $i$, yielding $z=26$. Furthermore, polymers are monodisperse with $N_{\text{m}}$ = 32, such that the volume occupied by a polymer chain is $v_{\text{p}}=32v_{0}$. The volume occupied by each solvent particle and cosolvent particle is $v_{0}$. Particle orientations are restricted to the 26 lattice directions. The Flory-Huggins-Potts framework can, in principle, double the number of interaction parameters relative to Flory–Huggins theory because each pair of species admits distinct aligned and misaligned energies. To keep the model tractable while still probing orientation effects, we introduce only three new parameters. Alignment-biased correlation network interactions are assigned between monomer–cosolvent and solvent–cosolvent pairs with fixed weights $p_{v}=1.0$ and $p_{\Omega}=0.25$ and the energetic penalty for misalignment is fixed $\Delta_{\text{mc}}=\Delta_{\text{sc}}=\Delta$, leaving all other contacts isotropic. This minimal extension is sufficient to capture the anisotropic mechanisms discussed below while avoiding an unwieldy parameter space. All simulations use a simple-cubic lattice with coordination number $z=26$ (nearest, next-nearest, and next-next-nearest neighbors). Energies are quoted in units of a fundamental scale $\epsilon_{0}=k_{\text{B}}T$.
At 300 K, representative non-covalent interactions span a wide energy range: water-to-amide hydrogen bonds are roughly –3 to –10 kcal/mol,68 $\pi-\pi$ contacts lie between 3-3- 3 and 5-5- 5 kcal/mol, 69, 70 whereas water–alcohol hydrogen bonds are much weaker, about 0.5-0.5- 0.5 kcal/mol.71, 72 To bracket this spectrum we scanned three anisotropic energies, $\Delta=-0.2\epsilon_{0},-0.5\epsilon_{0},-0.8\epsilon_{0}$.
@@ 2.3 Characterization of phase behavior
For select systems that meet the criterion above, their phase behavior is more precisely characterized by mapping binodal boundaries and determining the compositions of coexisting phases. Binodal boundaries are computed by first identifying critical points using constraints involving third-order derivatives 73, 74:
@@
Following identification of a critical point, coexistence lines are extended using a recently proposed natural parameter continuation algorithm75 akin to Gibbs-Duhem integration. As this approach iteratively identifies different compositions with equivalent chemical potentials, it also facilitates the construction of tie lines to indicate the equilibrium compositions of the coexisting phases. For additional details regarding this algorithm or the utilization of eqs. (8)-(10), readers are referred to Ref.  75 and the references therein.
@@ 2.4 Monte Carlo simulations
To characterize cononsolvency at a microscopic level, we perform lattice Monte Carlo (MC) simulations. The MC simulations directly implement specific manifestations of the Hamiltonian defined by eq. (1). These simulations are performed only for systems with parameters that exhibit mixture instability (Table 1). All simulations feature a $\times 34\times 34$ 34 simple cubic lattice with periodic boundaries and unit edge length; lattice sites are occupied by a mixture of solvent and cosolvent particles and a single polymer chain ($N_{\text{m}}=32$). For each manifestation of a Hamiltonian, simulations are performed over a set of cosolvent fractions that approximately (within the limits afforded by the discrete lattice) correspond to $x_{\text{c}}\in[0,1]$ in increments of 0.1.

### v3

@@
Department of Chemical and Biological Engineering, Princeton University, Princeton, NJ 08544
Cononsolvency occurs when mixing two good solvents creates poor-solvent conditions for polymers over specific composition ranges, causing macroscopic phase separation or microscopic chain collapse. Despite its technological and biophysical relevance, the connection between macroscopic and microscopic manifestations of cononsolvency remains unclear. A key challenge is identifying which interactions govern cononsolvency: coarse-grained analyses like standard Flory–Huggins models assume purely isotropic interactions, while atomistic simulations contain complex anisotropic interactions that cannot be precisely controlled or isolated. Here, we address the role of interaction anisotropy using the Flory-Huggins-Potts framework, which yields $\chi$ as a thermodynamic average over both configurational and internal-state coarse-grained degrees of freedom. This enables controlled comparison between systems with isotropic versus orientation-dependent interactions that share identical effective $\chi$ parameters, either driving cononsolvency by strong solvent-cosolvent affinity or preferential polymer-cosolvent affinity. While pairs of systems exhibit equivalent macroscopic phase behavior, lattice Monte Carlo simulations reveal that those featuring orientation-dependent interactions generate distinct collapse signatures, particularly in reentrant coil-globule transitions or characteristics of the solvation structure. These results demonstrate how microscopic interactions influences cononsolvency behavior beyond what effective $\chi$ parameters alone predict.
@@ 1 Introduction
Multiple mechanisms of polymer cononsolvency have been proposed on the basis of theoretical, computational, and experimental analyses. The preferential mixing mechanism attributes polymer collapse to favorable solvent–cosolvent interactions, where the polymer sacrifices conformational entropy to promote solvent–cosolvent mixing.13, 4, 16, 34, 35, 36 The seminal experimental study by Schild et al.37 demonstrated, however, that perturbing water–methanol interactions ($\chi_{12}$) is insufficient to predict cononsolvency of PNIPAM (poly(N-isopropylacrylamide)) in a methanol–water solution. An alternative mechanism involves preferential polymer–cosolvent adsorption, where trace amounts of cosolvent selectively solvate the polymer and form “enthalpic bridges” between distal chain segments.38 These interactions can trigger phase separation or single-chain collapse.39, 40, 41, 42, 43, 44, 14, 45, 46 Effects akin to preferential adsorption can also arise from entropic factors, such as size asymmetry between solvent and cosolvent leading to effective depletion interactions.36 In another scenario, the cosolvent can act like a surfactant, such that polymer contraction allows the cosolvent to interact favorably with both solvent and polymer.47 The concept of geometric frustration has also been used to explain cononsolvency, proposing that competition between cosolvent and solvent for polymer solvation destabilizes the local environment.48, 49 Overall, these studies highlight polymer cononsolvency as a phenomenon with various competing physicochemical interactions.
@@
Previous studies using Flory-Huggins (FH) theory have demonstrated that phase separation can be induced when either solvent-cosolvent or polymer-solvent interactions dominate, even when the polymer is miscible with each solvent individually.34, 50, 51 Furthermore, analyses of structure factors obtained following application of the random phase approximation provide similar conclusions. 52 Nevertheless, mean-field insights into the energetic factors driving phase separation may not always translate into understanding microscopic physics, such as single-chain conformational behavior.38, 53 In this direction, field theoretic methods14, 46 have demonstrated that single-chain collapse indeed occurs given strong, preferential affinity of the polymer for one of the solvents, but there is no apparent coil-globule transition driven purely by solvent-cosolvent affinity. Recent developments in variational field-theoretic treatments54 may enable further insights into single-chain behavior under such conditions.
@@
Molecular dynamics (MD) simulations of systems with generic chemical attributes (e.g., bead-spring polymers in monomeric solvents) have effectively illustrated many of the aforementioned mechanisms of cononsolvency at a microscopic level.36, 38, 55 A typical observation from such simulations is that the $R_{\text{g}}$ for a single polymer chain first decreases and then increases as cosolvent is progressively added. This scenario has been induced by solvent-cosolvent mixing, preferential polymer-solvent adsorption, and the surfactant-like action of the cosolvent by analyzing the molecular environment of the polymer and carefully evaluating molecular affinities. 38, 56, 57, 58, 49
MD simulations have also been used to examine the behavior of chemically specific systems, such as PNIPAM in water and ethanol. In such cases, single-chain collapse, consistent with cononsolvency, has been observed. Despite significant interest and detailed study, there is debate as to what precise molecular interactions or combination thereof begets this behavior.59, 60, 61, 62, 63 This highlights how microscopic drivers of cononsolvency can be non-trivial to resolve, even if the basic physics are well understood. This study does not adjudicate the specific atomistic interactions driving cononsolvency–it disentangles how systems with the same effective $\chi$ parameters show distinct molecular behavior.
@@
Several recent works have established quantitative connections between microscopic chain behavior and macroscopic phase separation in cononsolvency for systems with isotropic interactions. Zhang has comprehensively mapped the energetic conditions (i.e., $\chi$ parameters) that yield cononsolvency through mean-field analysis.51, 64 Previously, Zhang et al. also demonstrated via Wang-Landau simulations that trace cosolvent addition induces continuous coil-globule-coil transitions without requiring explicit solvent-solvent attraction.65 By a complmenetary and distinct approach, Marcato et al. mapped lattice models onto $\mathcal{O}(n)$-vector spin models to derive exact partition functions and field-theoretic descriptions amenable to analytics.66 Li et al. have established connections between single-chain collapse, multi-chain aggregation, and mean-field theory for block copolymers and homopolymer chain(s) in binary solvents, in the context of FH models67 While these studies establish that specific combinations of scalar $\chi$ parameters derived from isotropic interactions produce cononsolvency, this leaves unresolved whether orientation-dependent interactions fundamentally alters cononsolvency mechanisms or merely modulates existing ones.
@@
While simple FH theory captures cononsolvency phenomena, the apparent connection between thermoresponsive polymer solutions and cononsolvency41, 68, 2 and highlighted importance of hydrogen bonding,69, 41 ultimately suggest that orientation-dependent interactions may be relevant for at least some specific cononsolvency mechanisms. Here, we directly investigate how anisotropic interactions influences polymer cononsolvency in the context of a Flory-Huggins-Potts (FHP) framework,70 which previously demonstrated that including orientation-dependent interactions within FH-like models enabled description of thermoresponsive phenomena (e.g., miscibility loops and heating-induced coil-globule transitions) without temperature-dependent $\chi$ parameters. This work does not resolve mechanisms or debate for any specific system; instead, it clarifies how orientation-dependent interactions, like hydrogen bonding or even packing effects, influence cononsolvency phenomena within a single, tractable framework. In particular, we demonstrate how systems with the same effective $\chi$ interaction parameters, and thus identical macroscopic phase behavior, may exhibit different microscopic signatures of cononsolvency.
The remainder of the paper is organized as follows. After presentation of methods, we first identify established phase-separation regimes driven by solvent-cosolvent mixing versus polymer-cosolvent adsorption. Then, we compare corresponding coil-globule-coil transitions across systems with identical effective $\chi$ parameters but different underlying energetic contributions using lattice Monte Carlo simulations. Subsequent analysis reveals how anisotropic interactions generate microscopically distinct collapse pathways compared to isotropic systems, despite yielding identical macroscopic $\chi$ parameters. These findings provide additional insights into cononsolvency, by ascertaining the influence of additional degrees of freedom imparted by the FHP framework, and point towards strategies for tuning polymer response based on molecular-level interactions.
@@ 2.1 Flory-Huggins-Potts framework and stability analysis
To enable description of orientation-dependent interactions, we extend the recently developed FHP framework70 to model ternary polymer solutions. In effect, the FHP framework augments conventional FH theory by assigning internal states to particles that modulate pairwise interactions. These states may be interpreted, for example, as capturing aspects of molecular orientation at a coarse-grained resolution, enabling expression of orientation-dependent interactions, whereas traditional FH assumes isotropic interactions. While the motivation and detailed development of the FHP framework can be found in Ref. 70 , we briefly review some salient aspects.
@@
where $\Theta(\cdot)$ is the Heaviside step function, and $\delta$ defines the angular tolerance for alignment. We refer to this interaction style as a “correlation network,” and it resembles interactions in Maier-Saupe theory.71 Correlation networks allow all neighbors to align cooperatively, since the function becomes nonzero when two vectors point nominally in the same direction. Importantly, the orientation vectors here are coarse-grained internal state variables that modulate local interaction energies. In one manifestation, these vectors could encode differences in molecular orientation, where the positioning of certain functional groups would influence the manner of interaction with other moieties. However, these vectors may also represent other forms of collective ordering, such as that seen in nematic liquid crystals or in hydrogen-bond networks in water, where many simultaneous, orientation-dependent interactions occur. While $\Delta_{\alpha_{i}\alpha_{j}}$ terms are not treated in chemically specific terms here, surveying its influence over different values phenomenologically probes the role of orientation-dependent interactions, irrespective of their origin.
We note that the transition from simple FH to FHP introduces several additional parameters governing particle interactions. However, not all parameters are essential for modeling specific phenomena, and FHP parameters have been successfully fitted to experimental data with high precision.70 Regularization techniques and physical constraints can yield parsimonious models when needed. Since our objective is to distinguish cononsolvency manifestations, we employ minimal models that introduce vary only two while keeping other variables fixed. In particular, alignment-biased correlation network interactions are assigned between monomer–cosolvent and solvent–cosolvent pairs with fixed weights $p_{v}=1.0$ and $p_{\Omega}=0.25$ and the energetic penalty for misalignment is fixed $\Delta_{\text{mc}}=\Delta_{\text{sc}}=\Delta$; all other interactions are isotropic. This minimal extension is sufficient to capture the anisotropic mechanisms discussed below while avoiding an unwieldy parameter space. To provide some physical context, in the main text, we report results using $\Delta=0.5\epsilon_{0}$ where $\epsilon_{0}=k_{\text{B}}T$ is the unit of energy. At 300 K, representative non‑covalent interactions span a wide energy range: water‑to‑amide hydrogen bonds are roughly –3 to –10 kcal/mol,72 $\pi-\pi$ contacts lie between 3-3 and 5-5 kcal/mol, 73, 74 whereas water–alcohol hydrogen bonds are much weaker, about 0.5-0.5 kcal/mol.75, 76 To bracket this spectrum, additional results for $\Delta=-0.2\epsilon_{0},-0.5\epsilon_{0},-0.8\epsilon_{0}$ are provided in the Supplementary Information.
@@ 2.2 Helmholtz energy by mean-field analysis
Following Ref. 70 , but extended to three components, the intensive free energy of mixing per lattice site is:
@@
The terms $p_{v}$ and $p_{\Omega}$ are geometric factors related to what defines aligned interactions. In particular, $p_{v}$ is the fraction of neighbors that can form aligned interactions, and $p_{\Omega}$ is the fraction of pairwise orientations that are classified as aligned for such neighbors. While eq. 4 retains the classical FH functional form dictated by symmetry and thermodynamic consistency, eqs. 5 and 6 allow $\chi$-values to incorporate state-dependent interactions into the FH framework without modifying its structure. Thus, the key offering of leveraging the FHP conceptual framework is that the various terms that contribute to $\chi^{\text{FHP}}_{jk}$ are traceable to a well-defined and immutable microscopic Hamiltonian with controllable orientation-dependent energy terms.
For simplicity, all numerical results correspond to a simple cubic lattice where $\mathcal{N}(i)$ includes the nearest, next-nearest, and next-next-nearest neighbors of site $i$, yielding $z=26$. Furthermore, polymers are monodisperse with $N_{\text{m}}$ = 72, such that the volume occupied by a polymer chain is $v_{\text{p}}=N_{\text{m}}v_{0}$. The volume occupied by each solvent particle and cosolvent particle is $v_{0}$. Particle orientations are restricted to the 26 lattice directions. Although choices regarding the lattice and the interaction neighborhood influences $z$, a different prescription is expected to only modify the energy scales at which phenomena are observed without qualitatively altering mechanisms.
@@ 2.3 Characterization of phase behavior
$\mathbf{H}$ $=\begin{pmatrix}\frac{\partial^{2}\beta\Delta f_{\text{mix}}}{\partial\phi_{\text{p}}^{2}}&\frac{\partial^{2}\beta\Delta f_{\text{mix}}}{\partial\phi_{\text{p}}\partial\phi_{\text{s}}}\\ \frac{\partial^{2}\beta\Delta f_{\text{mix}}}{\partial\phi_{\text{p}}\partial\phi_{\text{s}}}&\frac{\partial^{2}\beta\Delta f_{\text{mix}}}{\partial\phi_{\text{s}}^{2}}\end{pmatrix}$ (8)
@@
For select systems that meet the criterion above, their phase behavior is more precisely characterized by mapping binodal boundaries and determining the compositions of coexisting phases. Binodal boundaries are computed by first identifying critical points using constraints involving third-order derivatives 77, 78:
@@
Following identification of a critical point, coexistence lines are extended using a recently proposed natural parameter continuation algorithm79 akin to Gibbs-Duhem integration. As this approach iteratively identifies different compositions with equivalent chemical potentials, it also facilitates the construction of tie lines to indicate the equilibrium compositions of the coexisting phases. For additional details regarding this algorithm or the utilization of eqs. (8)-(11), readers are referred to Ref. 79 and the references therein.
@@ 2.4 Monte Carlo simulations
To characterize cononsolvency at a microscopic level, we perform lattice Monte Carlo (MC) simulations. The MC simulations directly implement specific manifestations of the Hamiltonian defined by eq. (1). These simulations are performed only for systems with parameters that exhibit mixture instability (Table 1). All simulations feature a $34\times 34\times 34$ simple cubic lattice with periodic boundaries and unit edge length; lattice sites are occupied by a mixture of solvent and cosolvent particles and a single polymer chain ($N_{\text{m}}=32$). For each manifestation of a Hamiltonian, simulations are performed over a set of cosolvent fractions that approximately (within the limits afforded by the discrete lattice) correspond to $x_{\text{c}}\in[0,1]$ in increments of 0.1.
@@
It is worth nothing that the MC simulations technically reflect results for a chain in the canonical ensemble. Consequently, it is conceivable that substantial partitioning of solvent species into the near-space volume occupied by the polymer could reduce the effective (co)solvent mole fractions in the “bulk,” and such behavior would depend on the finite size of the simulation cell. Here, we verified that the number of solvent and cosolvent molecules coordinated within the polymer solvation shell remains small relative to the total particle count ($0.5\%$, see SI, Fig. S2), ensuring that bulk composition remains effectively constant, and this should be sufficient to effectively draw the connection between macroscopic thermodynamics and the dilute single-chain physics. In the future, it may be preferable to employ a constant chemical potential ensemble,14, 45 where the polymer gyration volume exchanges solvent and cosolvent with an external reservoir, though this may require additional considerations to be compatible with the Hamiltonians explored here.